OpenStudy (anonymous):

Given that 2sin^2 theta + sin theta -1 = 0, find the two values for sin theta.

OpenStudy (anonymous):

\[2 \sin ^{2} \theta + \sin \theta -1 = 0\, \] find the two values for \[\sin \theta\]

OpenStudy (kunal):

theta = n*pi

OpenStudy (kunal):

\[\sin \theta = (-1 + \sqrt{(-1)^2 -4(2)(-1)})/2\]

OpenStudy (kunal):

that is quadratic formula

OpenStudy (kunal):

\[\sin \theta = 1 => \theta = n *\pi\]

OpenStudy (anonymous):